The distance a basketball player runs during a game is one of the many variables used for assessing the performance of the players. 79 players are followed for a game each and the average distance ran by them is found to be 3457 meters (m) with a standard deviation of 147 m. Then the lower limit for the 95% confidence interval for the true mean is 


1
Expert's answer
2021-07-16T08:35:04-0400

The critical value for "\alpha=0.05" and "df=n-1=79-1=78"  degrees of freedom is "t_c=z_{1-\alpha/2, n-1}=1.990847."  

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(3457-1.990847\times\dfrac{147}{\sqrt{79}}, 3457+1.990847\times\dfrac{147}{\sqrt{79}})"

"=(3424.074, 3489.926)"

The lower limit for the 95% confidence interval for the true mean is 3424.074 m.